In the hydrogen atom, what is the electric potential energy of the electron when it is found in the n = 3 state? Hint: refer to the derivation of the total energy for energy proportions.
Answer #1

For an electron revolving in a circular orbit of radius, r around a nucleus with Z positive charge,

PE = -Ze2/r

KE = Ze2/2r

Hence:

Etotal = (-Ze2/r) + (Ze2/2r) = -Ze2/2r

and

energy of bohr's orbit

Since Z = 1 for hydrogen above equation can be further simplified to:

En = -13.6/n2 eV

so, potential energy is double of total energy with negative sign

En = -2x13.6/n2 eV

so, for n = 3

E3 = - 2x13.6/3^2 eV = -3.02 eV.............................ANS